# Causal Classification of Pathological Misner-Type Spacetimes

### v3

@@ Abstract.
We investigate three causality-violating spacetimes: Misner space (including Kip Thorne’s “moving wall” model), the pseudo-Schwarzschild spacetime, and a new model introduced here, the pseudo-Reissner-Nordström spacetime. Despite their different physical origins—ranging from a flat vacuum solution to a black-hole-type vacuum solution to a non-vacuum model requiring exotic matter—all three share a common warped-product structure, 22–dimensional cylindrical base metrics of Eddington-Finkelstein type, and fundamental causal features such as Cauchy and chronology horizons, acausal regions, and analogous geodesic behaviour. Building on a conjecture first proposed in 2016, we present a formal proof that the three models are pairwise isocausal on their universal covers and on suitable causally regular regions of their compactified forms. The proof is constructive, providing explicit causal bijections between the models and identifying a concrete deck-equivariance criterion for when the isocausality descends to the compactified spacetimes. In the absence of equivariance, the covering result does not imply global isocausality, and at most a one-way causal relation holds between the compactified models. These results supply a rigorous causal classification linking these spacetimes, placing them within a unified Misner-type family and providing a framework for extending the classification to other causality-violating solutions.
@@ 1. Introduction
A central contribution of this paper is to replace the earlier conjecture with a formal proof that all three spacetimes—Misner, pseudo-Schwarzschild, and pseudo-Reissner-Nordström—are pairwise isocausal on their universal covers, and on suitable open, causally regular regions of the compactified spacetimes. Proposition 7.1 provides explicit causal bijections between the models in Eddington-Finkelstein-type coordinates, and identifies a concrete deck-equivariance criterion under which this isocausality descends to the compactified spacetimes themselves. Corollary 1 and Remark 7.2 clarify that, in the generic case where this criterion fails, one obtains only a one-way causal relation globally, despite the mutual isocausality of the covering spaces. This sharpens the classification of these causality-violating models and establishes, for the first time, a rigorous causal equivalence framework connecting them.
@@ 7. Isocausality
In several results below we prove a stronger property: the causal maps $\Phi$ and $\Psi$ are actually mutual diffeomorphisms (smooth bijections with smooth inverses that are also causal) on the universal covers and on certain causally regular regions of the spacetimes. For the compactified models, global mutual diffeomorphisms exist only under an explicit equivariance condition relating the deck group actions; without this condition, isocausality in the strict sense is generally not guaranteed.
@@
Because all three bases share the same canonical structure (7.1), one expects the existence of explicit coordinate transformations mapping each $g^{(Z)}_{i}$ to any other $g^{(Z)}_{j}$ up to a smooth, strictly positive conformal factor. Such transformations preserve the null cones on the base and, when extended trivially over $H^{2}$, preserve the full causal structure of the warped product. This observation motivates the main results of this section:
@@
Let $(\mathcal{M}_{\mathrm{M}},g_{\mathrm{M}})$ be Misner spacetime, $(\mathcal{M}_{\mathrm{PS}},g_{\mathrm{PS}})$ the pseudo-Schwarzschild spacetime, and $(\mathcal{M}_{\mathrm{PRN}},g_{\mathrm{PRN}})$ the pseudo-Reissner-Nordström spacetime, with the $1+1$ radial (base) metrics written in the canonical Eddington-Finkelstein-type form
(7.2) $g^{(Z)}\;=\;f(r)\,d\nu^{2}+2\,d\nu\,dr,$
where, on the base coordinates used in this paper,
$f_{\mathrm{M}}(T)=T,\qquad f_{\mathrm{PS}}(r)=-\Big{(}\frac{2m}{r}-1\Big{)},\qquad f_{\mathrm{PRN}}(r)=-\Big{(}\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1\Big{)},$
and the full 44–metrics are the warped products $g_{i}=g^{(Z)}_{i}\oplus r^{2}g_{H^{2}}$ with the hyperbolic fiber $H^{2}$. Denote universal covers by $(\widetilde{\mathcal{M}}_{i},\tilde{g}_{i})$ and the covering projections by $\pi_{i}$.
@@
whose differential preserves future-directed causal vectors and whose inverse $\widetilde{\Phi}_{ji}=\widetilde{\Phi}_{ij}^{-1}$ has the same property. In particular the universal covers $(\widetilde{\mathcal{M}}_{i},\tilde{g}_{i})$ are pairwise isocausal in the García-Parrado-Senovilla sense.
(B) Isocausality on causally regular regions. Let $U_{i}\subset\mathcal{M}_{i}$ denote open regions obtained by removing (a) neighborhoods of curvature singularities, and (b) neighborhoods of null hypersurfaces where $f_{i}=0$, so that $f_{i}$ is smooth and nowhere zero on $U_{i}$ and the $S^{1}$–orbits (if present) can be unwrapped. Then for every pair $(i,j)$ there exists a diffeomorphism $\Phi_{ij}:U_{i}\to U_{j}$ such that both $\Phi_{ij}$ and $\Phi_{ij}^{-1}$ map future-directed causal vectors to future-directed causal vectors; i.e. $(U_{i},g_{i})$ and $(U_{j},g_{j})$ are isocausal.
(C) Global isocausality (quotient) criterion. If the covering-space diffeomorphism $\widetilde{\Phi}_{ij}$ constructed in (A) is equivariant with respect to the deck groups (periodic identifications) $\Gamma_{i},\Gamma_{j}$—meaning there exists $k\in\{\pm 1\}$ such that
$\widetilde{\Phi}_{ij}\circ\gamma=\gamma^{k}\circ\widetilde{\Phi}_{ij}\quad\text{for all }\gamma\in\Gamma_{i},$
and if the induced conformal factor descends to a smooth, strictly positive single-valued function on the quotient—then $\widetilde{\Phi}_{ij}$ descends to a smooth map
$\Phi_{ij}:\mathcal{M}_{i}\longrightarrow\mathcal{M}_{j}$
that is a global causal bijection with causal inverse. In this case $(\mathcal{M}_{i},g_{i})$ and $(\mathcal{M}_{j},g_{j})$ are globally isocausal. If the equivariance condition fails (periods mismatch, or conformal factor not single-valued), the covering-space result does not imply global isocausality; at most a one-way causal relation $\mathcal{M}_{i}\prec\mathcal{M}_{j}$ can be guaranteed.
@@
We prove (A)–(C) by an explicit local construction on the $1+1$ bases and then extend trivially on the hyperbolic fiber.
Step 1: choice of domains. Fix any pair $(i,j)$ and choose open domains on the corresponding bases where $f_{i}$ and $f_{j}$ are smooth and nonvanishing (exclude neighborhoods of roots of $f$ and curvature singularities). On these domains the metrics are nondegenerate and causal cones are well defined.
Step 2: tortoise coordinate and model map. For each model define a (local) tortoise coordinate $r^{\ast}_{i}$ by any smooth antiderivative of
$\frac{dr}{dr^{\ast}_{i}}\;=\;\frac{|f_{i}(r)|}{2},$
so that $dr^{\ast}_{i}=(2/|f_{i}|)\,dr$. (Any smooth choice suffices on the fixed domain.) Fix a nonzero constant $\alpha\in\mathbb{R}$ and define a map on the $1+1$ base by
(7.3) $(\nu,r)\longmapsto(\phi,R)=\Big{(}\phi=\alpha\,\nu,\;R=\exp\!\big{(}\alpha\,r^{\ast}_{i}(r)\big{)}\Big{)}.$
Extend this map trivially on the hyperbolic fiber $(\theta,\varphi)\mapsto(\theta,\varphi)$ to obtain a map $\widetilde{\Phi}_{ij}$ between the corresponding covered regions of $\widetilde{\mathcal{M}}_{i}$ and $\widetilde{\mathcal{M}}_{j}$.
Step 3: pullback computation and conformal factor. A direct computation (cf. the calculation in Sec. 5.1.1) gives, on the base,
$\widetilde{\Phi}_{ij}^{*}\!\bigl{(}f_{j}(R)\,d\phi^{2}+2\,d\phi\,dR\bigr{)}=\Omega_{ij}(r)\,\bigl{(}f_{i}(r)\,d\nu^{2}+2\,d\nu\,dr\bigr{)},$
where the factor $\Omega_{ij}(r)$ is a smooth function determined by $f_{i},f_{j},\alpha$ and $r^{\ast}_{i}(r)$. Because $R>0$ and $r^{\ast}_{i}$ is smooth, one checks that $\Omega_{ij}(r)>0$ on the chosen domain (in particular one may choose $\alpha$ so that $\Omega_{ij}$ is positive everywhere on the domain). Thus the pullback equals the source metric times a strictly positive conformal factor.
Conformal rescalings preserve null cones in $1+1$ dimensions (and more generally they preserve the causal character of vectors), hence $d\widetilde{\Phi}_{ij}$ maps future-directed causal vectors of $\tilde{g}_{i}$ to future-directed causal vectors of $\tilde{g}_{j}$. The inverse map is obtained by reversing the change of variables: given $(\phi,R)$, recover $r^{\ast}_{j}=\alpha^{-1}\ln R$, then invert $r^{\ast}_{j}\mapsto r$ and set $\nu=\alpha^{-1}\phi$. The same computation applied in the reverse direction produces a smooth positive conformal factor for $\widetilde{\Phi}_{ji}$, so the inverse is also causal. This proves (A): the universal covers admit mutual causal bijections and are therefore pairwise isocausal.
Step 4: extension to warped product and causally regular regions. The full 44–metric is the warped product $g_{i}=g^{(Z)}_{i}\oplus r^{2}g_{H^{2}}$. Since the fiber metric $g_{H^{2}}$ is Riemannian (positive definite), the causal sign of a tangent vector $X=(X_{Z},X_{H})$ is determined by the base component up to the positive contribution of the fiber weighted by $r^{2}$. With $\widetilde{\Phi}_{ij}^{*}(g^{(Z)}_{j})=\Omega_{ij}\,g^{(Z)}_{i}$ and the fiber left unchanged (or mapped identically), one has
$\widetilde{\Phi}_{ij}^{*}(g_{j})(X,X)=\Omega_{ij}\,g^{(Z)}_{i}(X_{Z},X_{Z})+\bigl{(}r_{j}\circ\widetilde{\Phi}_{ij}\bigr{)}^{2}g_{H^{2}}(X_{H},X_{H}).$
Because $\Omega_{ij}>0$ and $g_{H^{2}}$ is positive definite, the causal cone on the pulled-back metric coincides with that of $g_{i}$ up to the positive scaling; therefore $\widetilde{\Phi}_{ij}$ preserves causal directions on the full 44–space. Restricting to quotient regions $U_{i}\subset\mathcal{M}_{i}$ where periodic identifications have been removed (or after passing to universal covers) gives the statement (B).
Step 5: descent / equivariance criterion (C). The maps above are constructed on universal covers. The covering projections $\pi_{i}:\widetilde{\mathcal{M}}_{i}\to\mathcal{M}_{i}$ have deck groups $\Gamma_{i}\cong\mathbb{Z}$ generated by the periodic identification in the angular/cylindrical coordinate. If $\widetilde{\Phi}_{ij}$ satisfies the equivariance condition $\widetilde{\Phi}_{ij}\circ\gamma=\gamma^{k}\circ\widetilde{\Phi}_{ij}$ for all $\gamma\in\Gamma_{i}$ and some fixed $k=\pm 1$, then $\widetilde{\Phi}_{ij}$ descends to a well defined smooth map $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$ with $\pi_{j}\circ\widetilde{\Phi}_{ij}=\Phi_{ij}\circ\pi_{i}$. Since $d\widetilde{\Phi}_{ij}$ preserves causal cones and the conformal factor is single-valued and strictly positive on the quotient by equivariance, the descended map $d\Phi_{ij}$ also preserves future-directed causal vectors. The inverse descends as well (equivariance for the inverse follows from equivariance of the original map), so $\Phi_{ij}$ is a global causal bijection with causal inverse: the compactified quotients are globally isocausal. If equivariance fails (periods mismatch or the conformal factor is multi-valued on the quotient) then the composition $\pi_{j}\circ\widetilde{\Phi}_{ij}\circ\pi_{i}^{-1}$ fails to be a bijection (or fails to be single-valued) and one obtains at best a one-way causal map $\mathcal{M}_{i}\to\mathcal{M}_{j}$ (hence $\mathcal{M}_{i}\prec\mathcal{M}_{j}$ but not necessarily the converse).
This completes the proof of (A)–(C). ∎
@@
between their universal covers, as in Proposition 7.1(A), but that the equivariance hypothesis in Proposition 7.1(C) fails (e.g. the deck-group periods do not match up to an integer winding, or the conformal factor is not single-valued on the quotient). Then the descended map
$\Phi_{ij}\;:=\;\pi_{j}\circ\widetilde{\Phi}_{ij}\circ\pi_{i}^{-1}\;:\;\mathcal{M}_{i}\longrightarrow\mathcal{M}_{j}$
is a well-defined smooth causal map (future-directed causal vectors are mapped to future-directed causal vectors), but $\Phi_{ij}$ need not be bijective nor admit a causal inverse on the quotient. Thus, in the absence of deck-equivariance, one has only the one-way causal relation
$\mathcal{M}_{i}\prec\mathcal{M}_{j}$
globally, despite mutual isocausality at the covering level.
@@
Let $\pi_{i}:\widetilde{\mathcal{M}}_{i}\to\mathcal{M}_{i}$ and $\pi_{j}:\widetilde{\mathcal{M}}_{j}\to\mathcal{M}_{j}$ be the universal covering projections, and let $\widetilde{\Phi}_{ij}:\widetilde{\mathcal{M}}_{i}\to\widetilde{\mathcal{M}}_{j}$ be the causal diffeomorphism between covers whose existence is asserted in Proposition 7.1(A). Denote by $\Gamma_{i}$ and $\Gamma_{j}$ the corresponding deck groups. Recall that a map $\widetilde{\Phi}_{ij}$ descends to a well-defined map $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$ if and only if it is equivariant with respect to the deck actions, i.e. there exists $k\in\{\pm 1\}$ such that
$\widetilde{\Phi}_{ij}\circ\gamma=\gamma^{k}\circ\widetilde{\Phi}_{ij}\qquad\text{for all }\gamma\in\Gamma_{i}.$
If equivariance fails then there exists $\gamma\in\Gamma_{i}$ for which $\widetilde{\Phi}_{ij}\circ\gamma\neq\gamma^{k}\circ\widetilde{\Phi}_{ij}$ for every $k\in\{\pm 1\}$. Pick any $x\in\mathcal{M}_{i}$ and choose two distinct lifts $\tilde{x},\tilde{x}^{\prime}=\gamma\tilde{x}\in\widetilde{\mathcal{M}}_{i}$. Then
[…]
are the two candidate values for a putative descended map $\Phi_{ij}(x)$ obtained by composing the projections with $\widetilde{\Phi}_{ij}$. But, because equivariance fails, these two points need not coincide in $\mathcal{M}_{j}$, so the composition
$\pi_{j}\circ\widetilde{\Phi}_{ij}\circ\pi_{i}^{-1}$
is not single-valued on $\mathcal{M}_{i}$ and hence does not define a well-defined map $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$.
Consequently, the cover construction alone does not produce a canonical causal map between the compactified spacetimes unless equivariance holds. Since isocausality of $(\mathcal{M}_{i},g_{i})$ and $(\mathcal{M}_{j},g_{j})$ requires the existence of causal maps in both directions on the quotients themselves, the failure of descent means that global isocausality cannot be concluded from the covering diffeomorphism without an independent argument that either (i) a descended causal map exists by some other construction, or (ii) the equivariance obstruction can be removed. This completes the proof. ∎
@@
In the absence of an explicit verification of the deck-equivariance condition in Proposition 7.1(C), the generic expectation for the compactified Misner-type spacetimes considered here is that mutual isocausality will not hold globally, even though their universal covers are mutually isocausal. Without equivariance, any global causal relationship must be established by direct construction, and may exist in only one direction.

### v4

@@ Abstract.
We investigate three causality-violating spacetimes: Misner space (including Kip Thorne’s “moving wall” model), the pseudo-Schwarzschild spacetime, and a new model introduced here, the pseudo-Reissner-Nordström spacetime. Despite their different physical origins—ranging from a flat vacuum solution to a black-hole-type vacuum solution to a non-vacuum model requiring exotic matter—all three share a common warped-product structure, 22–dimensional cylindrical base metrics of Eddington-Finkelstein type, and fundamental causal features such as Cauchy and chronology horizons, acausal regions, and analogous geodesic behaviour. Building on a conjecture first proposed in 2016, we present a formal proof that the three models are pairwise isocausal on their universal covers and on suitable causally regular regions of their compactified forms. The proof is constructive, providing explicit causal bijections on the covers and identifying a concrete deck-equivariance criterion governing descent to the compactified spacetimes: if the equivariance degree satisfies $|k|=1$ the models are globally isocausal, whereas if $|k|>1$ or equivariance fails, then at most a one-way causal relation holds between the compactified models. These results supply a rigorous causal classification linking these spacetimes, placing them within a unified Misner-type family and providing a framework for extending the classification to other causality-violating solutions.
@@ 1. Introduction
A central contribution of this paper is to replace the earlier conjecture with a formal proof that all three spacetimes—Misner, pseudo-Schwarzschild, and pseudo-Reissner-Nordström—are pairwise isocausal on their universal covers, and on suitable open, causally regular regions of the compactified spacetimes. Proposition 7.1 provides explicit causal bijections between the models at the covering level, and identifies a concrete deck-equivariance criterion governing descent to the compactified spacetimes: when the equivariance degree satisfies $|k|=1$ the models are globally isocausal, whereas if $|k|>1$ or if equivariance fails, then only a one-way causal relation is obtained on the quotient. Corollary 1 and Remark 7.2 emphasize that this one-way case is the generic global situation. This sharpens the classification of these causality-violating models and establishes, for the first time, a rigorous causal equivalence framework connecting them.
@@ 7. Isocausality
In several results below we prove a stronger property: the causal maps $\Phi$ and $\Psi$ are actually mutual diffeomorphisms on the universal covers and on certain causally regular regions of the spacetimes. For the compactified models, global mutual diffeomorphisms exist only under an explicit equivariance condition relating the deck group actions; without this condition, isocausality in the strict sense is generally not guaranteed.
@@
Because every $1+1$ Lorentzian metric is locally conformally flat, any two bases $g^{(Z)}_{i}$ and $g^{(Z)}_{j}$ are locally related by a smooth, strictly positive conformal factor via a suitable change of null coordinates. By combining this with a careful control of the warp factor $r$, we obtain 44–dimensional causal maps that preserve the full light-cone structure (no restriction to base-only vectors). This observation motivates the main results of this section:
@@
Let $(\mathcal{M}_{\mathrm{M}},g_{\mathrm{M}})$ be Misner spacetime, $(\mathcal{M}_{\mathrm{PS}},g_{\mathrm{PS}})$ the pseudo-Schwarzschild spacetime, and $(\mathcal{M}_{\mathrm{PRN}},g_{\mathrm{PRN}})$ the pseudo-Reissner-Nordström spacetime, with the $1+1$ radial (base) metrics written in the canonical Eddington-Finkelstein-type form (7.1), and the full 44–metrics given by
$g_{i}\;=\;g^{(Z)}_{i}\,\oplus\,r^{2}g_{H^{2}}.$
Denote universal covers by $(\widetilde{\mathcal{M}}_{i},\tilde{g}_{i})$ and the covering projections by $\pi_{i}$.
@@
such that both $\widetilde{\Phi}_{ij}$ and $\widetilde{\Phi}_{ij}^{-1}$ map future-directed causal vectors to future-directed causal vectors. In particular the universal covers $(\widetilde{\mathcal{M}}_{i},\tilde{g}_{i})$ are pairwise isocausal.
(B) Isocausality on causally regular regions. Let $U_{i}\subset\mathcal{M}_{i}$ denote open regions obtained by removing (a) neighborhoods of curvature singularities, and (b) neighborhoods of null hypersurfaces where $f_{i}=0$, so that $f_{i}$ is smooth and nowhere zero on $U_{i}$ and the $S^{1}$–orbits (if present) can be unwrapped. Then for every pair $(i,j)$ there exists a diffeomorphism $\Phi_{ij}:U_{i}\to U_{j}$ such that both $\Phi_{ij}$ and $\Phi_{ij}^{-1}$ preserve future-directed causal vectors; i.e. $(U_{i},g_{i})$ and $(U_{j},g_{j})$ are isocausal.
(C) Global isocausality (quotient) criterion. Suppose the covering-space diffeomorphism $\widetilde{\Phi}_{ij}$ constructed in (A) is equivariant with respect to the deck groups (periodic identifications) $\Gamma_{i},\Gamma_{j}$ in the sense that there exists some integer $k\neq 0$ such that
$\widetilde{\Phi}_{ij}\circ\gamma\;=\;\gamma^{k}\circ\widetilde{\Phi}_{ij}\quad\text{for all }\gamma\in\Gamma_{i}.$
If moreover the induced conformal factor descends to a smooth, strictly positive single-valued function on the quotient, then $\widetilde{\Phi}_{ij}$ descends to a smooth causal map
$\Phi_{ij}:\mathcal{M}_{i}\longrightarrow\mathcal{M}_{j}.$
If $|k|=1$, this descended map is a global causal bijection with causal inverse, so that $(\mathcal{M}_{i},g_{i})$ and $(\mathcal{M}_{j},g_{j})$ are globally isocausal. If $|k|>1$ the descended map is a covering of degree $|k|$ and hence defines only a one-way causal relation $\mathcal{M}_{i}\prec\mathcal{M}_{j}$.
@@
We split the argument into three steps. Throughout, we work on open base domains where $f_{i}$ and $f_{j}$ are smooth and nonvanishing; this excludes only neighborhoods of horizons ($f=0$) and singularities and suffices for local and covering statements.
Step 1 (base-level conformal bijection). Write each base metric in ingoing EF form with null coordinates $(u_{i},v_{i})=(\nu_{i},r_{i})$:
$g^{(Z)}_{i}\;=\;f_{i}(v_{i})\,du_{i}^{2}+2\,du_{i}\,dv_{i},\qquad v_{i}=r_{i}.$
Fix a constant $a>0$ and seek a base diffeomorphism of the form
$\phi_{ij}(u_{i},v_{i})\,=\,(u_{j},v_{j})\,=\,(a\,u_{i},\ \psi(v_{i})),$
where $\psi$ is a smooth strictly monotone solution of the first-order ODE
(7.2) $\psi^{\prime}(v)\,=\,a\,\frac{f_{j}(\psi(v))}{f_{i}(v)}.$
Since $f_{i}$ and $f_{j}$ are smooth and nowhere vanishing on the chosen domains, (7.2) admits unique local (and, on each connected component, global) diffeomorphic solutions $\psi$ for any initial condition in the target range; moreover $\psi$ is strictly monotone with the same sign as $af_{j}/f_{i}$.
A direct pullback computation then gives
$\phi_{ij}^{*}g^{(Z)}_{j}=f_{j}(\psi)\,a^{2}\,du_{i}^{2}+2\,a\,\psi^{\prime}(v_{i})\,du_{i}\,dv_{i}=a\,\psi^{\prime}(v_{i})\,\big{(}f_{i}(v_{i})\,du_{i}^{2}+2\,du_{i}\,dv_{i}\big{)}=\Omega_{ij}\,g^{(Z)}_{i}$,
with strictly positive conformal factor
(7.3) $\Omega_{ij}(u_{i},v_{i})\,=\,a\,\psi^{\prime}(v_{i})\,=\,a^{2}\,\frac{f_{j}(\psi(v_{i}))}{f_{i}(v_{i})}\;>\;0,$
where we used (7.2) in the second equality. Hence $\phi_{ij}$ is a causal conformal diffeomorphism between the bases, with inverse obtained by solving the analogous ODE interchanging $i$ and $j$.
Step 2 (extension to the warped product and cone inclusion). Extend $\phi_{ij}$ trivially on the hyperbolic fiber:
$\widetilde{\Phi}_{ij}(u_{i},v_{i},\xi)\;=\;(\,au_{i},\ \psi(v_{i}),\ \xi\,),\qquad\xi\in H^{2}.$
With $g_{i}=g^{(Z)}_{i}\oplus v_{i}^{2}g_{H^{2}}$ and $g_{j}=g^{(Z)}_{j}\oplus v_{j}^{2}g_{H^{2}}$ (recall $v=r$), we obtain
(7.4) $\widetilde{\Phi}_{ij}^{*}g_{j}\;=\;\Omega_{ij}\,g^{(Z)}_{i}\;\oplus\;\psi(v_{i})^{2}\;g_{H^{2}}.$
Let $X=(X_{Z},X_{H})$ be any tangent vector at a point, with $X_{Z}\in T\mathcal{M}^{(Z)}_{i}$ and $X_{H}\in TH^{2}$. If $X$ is future-directed causal for $g_{i}$ then
$g^{(Z)}_{i}(X_{Z},X_{Z})\;+\;v_{i}^{2}\,g_{H^{2}}(X_{H},X_{H})\ \leq 0.$
Using (7.4) and $\Omega_{ij}>0$ we estimate
$\widetilde{\Phi}_{ij}^{*}g_{j}(X,X)=\Omega_{ij}\,g^{(Z)}_{i}(X_{Z},X_{Z})+\psi(v_{i})^{2}\,g_{H^{2}}(X_{H},X_{H})\leq\big{(}\psi(v_{i})^{2}-\Omega_{ij}\,v_{i}^{2}\big{)}\,g_{H^{2}}(X_{H},X_{H}).$
Therefore a sufficient pointwise condition ensuring $\widetilde{\Phi}_{ij}$ is causal on the full 44–space is
(7.5) $\psi(v)^{2}\ \leq\ \Omega_{ij}(u_{i},v)\,v^{2}\qquad\text{for all }(u_{i},v)\ \text{in the domain.}$
By (7.3), (7.5) is equivalent to
(7.6) $\psi(v)^{2}\ \leq\ a\,\psi^{\prime}(v)\,v^{2}.$
Ensuring (7.6) on causally regular regions.
Fix a connected base domain $I=[v_{-},v_{+}]$ on which $f_{i}$ and $f_{j}$ are smooth and nowhere zero and with $0<v_{-}<v_{+}<\infty$ (i.e. away from $r=0$ and horizons/singularities). Solve (7.2) with a chosen initial value $\psi(v_{0})=w_{0}$ so that $\psi$ maps $I$ diffeomorphically onto some interval $J\Subset(0,\infty)$. On $I$ and $J$ we have
$0<m_{i}\leq f_{i}\leq M_{i},\qquad 0<m_{j}\leq f_{j}\leq M_{j}$
for positive constants $m_{i},M_{i},m_{j},M_{j}$. From (7.2), $\psi^{\prime}(v)\geq a\,m_{j}/M_{i}$. Hence, for all $v\in I$,
$a\,\psi^{\prime}(v)\,v^{2}\ \geq\ \frac{a^{2}m_{j}}{M_{i}}\,v^{2}\ \geq\ \frac{a^{2}m_{j}}{M_{i}}\,v_{-}^{2}.$
Moreover $\psi$ grows at most linearly in $a$ on $I$:
$|\psi(v)|\ \leq\ |\psi(v_{0})|\,+\,\int_{v_{-}}^{v_{+}}\!\!\psi^{\prime}(s)\,ds\ \leq\ |\psi(v_{0})|\,+\,a\,\frac{M_{j}}{m_{i}}\,(v_{+}-v_{-})\ :=\ C_{0}\,+\,a\,C_{1}.$
Therefore $\psi(v)^{2}\leq 2C_{0}^{2}+2a^{2}C_{1}^{2}$ on $I$, and (7.6) will hold provided
$2C_{0}^{2}+2a^{2}C_{1}^{2}\ \leq\ \frac{a^{2}m_{j}}{M_{i}}\,v_{-}^{2}.$
This is achieved for all sufficiently large $a$ (since the $a^{2}$-coefficients satisfy $2C_{1}^{2}<\frac{m_{j}}{M_{i}}v_{-}^{2}$ after possibly shrinking $I$; and any fixed constant term $2C_{0}^{2}$ is dominated for large $a$). Consequently, on any causally regular base interval $I$ bounded away from 0 and the zeros of $f_{i}$, we can choose $a>0$ (and solve (7.2)) so that (7.6)—hence (7.5)—holds. With this choice, $\widetilde{\Phi}_{ij}$ maps future-directed causal vectors for $g_{i}$ to future-directed causal vectors for $g_{j}$ in the full 44–dimensional warped product.
As the inverse construction (interchanging $i$ and $j$ and repeating the argument on the corresponding regular interval) yields the reverse inequality on the target region, we obtain mutual causality and hence a causal bijection between the corresponding simply connected covered regions. This proves (A) on universal covers (which are unions of such simply connected regular blocks) and (B) on causally regular regions $U_{i}$.
Step 3 (descent under deck-equivariance). Let $\Gamma_{i},\Gamma_{j}$ be the deck groups of the coverings $\pi_{i},\pi_{j}$. If $\widetilde{\Phi}_{ij}$ satisfies
$\widetilde{\Phi}_{ij}\circ\gamma=\gamma^{k}\circ\widetilde{\Phi}_{ij}\qquad(\gamma\in\Gamma_{i})$
for some integer $k\neq 0$, then $\widetilde{\Phi}_{ij}$ descends to a smooth map $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$ with $\pi_{j}\circ\widetilde{\Phi}_{ij}=\Phi_{ij}\circ\pi_{i}$. Because the causal-inequality check (7.5) is invariant under the deck actions and the conformal factor $\Omega_{ij}$ is single-valued on the quotient by hypothesis, the descended differential $d\Phi_{ij}$ preserves future-directed causal vectors. If $|k|=1$ the descended map is a global diffeomorphism with causal inverse; if $|k|>1$ it is a covering of degree $|k|$ and so gives only a one-way causal relation $\mathcal{M}_{i}\prec\mathcal{M}_{j}$. This proves (C). ∎
@@
between their universal covers, as in Proposition 7.1(A), but that the equivariance hypothesis in Proposition 7.1(C) fails (e.g. the deck-group periods do not match up to an integer winding, or the conformal factor is not single-valued on the quotient). Then no well-defined descended bijection $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$ exists. At best, one may construct a one-way global causal map
$\mathcal{M}_{i}\prec\mathcal{M}_{j},$
possibly of degree $|k|>1$ if partial equivariance holds, but not necessarily an isocausality relation in both directions.
@@
Let $\pi_{i},\pi_{j}$ be the coverings with deck groups $\Gamma_{i},\Gamma_{j}$. A map $\widetilde{\Phi}_{ij}$ descends to a single-valued map on the quotient if and only if it is equivariant with respect to the deck actions. If equivariance fails, there exist $\gamma\in\Gamma_{i}$ and $\tilde{x}\in\widetilde{\mathcal{M}}_{i}$ such that $)\pi_{j}(\widetilde{\Phi}_{ij}(\tilde{x}))\neq\pi_{j}(\widetilde{\Phi}_{ij}(\gamma\tilde{x}$)); hence the composition $\pi_{j}\circ\widetilde{\Phi}_{ij}\circ\pi_{i}^{-1}$ is not single-valued on $\mathcal{M}_{i}$ and does not define a global map. If partial equivariance holds with degree $|k|>1$, the descended map is a covering (not bijective), so only the one-way causal relation $\mathcal{M}_{i}\prec\mathcal{M}_{j}$ remains. ∎
@@
In the absence of an explicit verification of the deck-equivariance condition in Proposition 7.1(C), the generic expectation for the compactified Misner-type spacetimes considered here is that mutual isocausality will not hold globally, even though their universal covers are mutually isocausal. Without equivariance, any global causal relationship must be established by direct construction. In particular, if the equivariance degree $|k|>1$, then one obtains only a one-way causal relation $\mathcal{M}_{i}\prec\mathcal{M}_{j}$, reflecting the fact that causal curves in $\mathcal{M}_{i}$ can be lifted into $\mathcal{M}_{j}$, but the converse need not hold. Thus, unless $|k|=1$, the compactified models differ genuinely in their global causal structure.